Analysis Shift Meeting Agenda, November 07 - November 13, 2021

From GlueXWiki
Jump to: navigation, search

Daily Analysis meeting at 2pm EST (CC room 326) BlueJeans Link

Saturday, November 13

  1. Projections
  2. Analysis Shift
  3. Timing of production runs
    1. What is raw mode and does that count in prod?
  4. Calibration Discussion
    1. Thanks, Sean!

Friday, November 12

  1. FCAL Calibration
    1. Igal will add calibrated plots here
  2. Updates
    1. TAGH plugin (Bhesha)
    2. TAGH and TAGM calibration (Bo and Bhesha)
    3. Plot Browser (Bo)
    4. Data Quality (Phoebe)
    5. Physics Analysis (Nathaly and Jackson)
  3. Target Scheduling Discussion
  4. Analysis Plan for the weekend
    1. Collect stats and work on calibrations
    2. Data monitoring
    3. Contingencies on target schedule
    4. Find flux

Thursday, November 11

  1. Updates
    1. Analysis Agenda!
    2. Analysis Shift Page
    3. Worker Shift Spreadsheet
    4. Getting .evio from jcache
      1. Igal needs production .evio files from
      2. Will be pinned for 60 days
  2. Questions for Sasha
    1. PS trigger rate measurement (where is it, what exists)
      1. Estimate trigger rate by PS production rate, use total triggers, crude estimate
      2. Takes a few runs, scales depending on .evio files
      3. Total trigger rate (kHz), PS trigger rate (kHz),
      4. How do you getting total triggers in a run?
        1. Process whole run, total PS triggers in tagger (H or M), and has to be calibrated
      5. Monitoring
        1. Compare 1 run vs another
        2. Look on specific channel
        3. sample you process, know how many triggers we have
        4. to get rel. yield/ flux
        5. Compare to another fraction of a run.
        6. 1st run you have 100 meson/ million triggers vs. 200/M then statistical errors, yield changes or stays the same.
        7. Usually 5 files/ run
    2. RF/TPOL
      1. Top Plot: Compare Rate changes between runs between detectors. General health of detectors
      2. Bottom plots: Make sure all channels are filled
    3. TPOL Intergral
    4. Trigger Run 64 &67
      1. Holes mean LED doesn't work. May need missing trigger is a good check on the BCAL energy
      2. Link to log entry about resetting pedestal [1]
  3. Processing Data
    1. .hddm is faster to process than .evio, but timing isn't calibrated correctly with hddm file
  4. Reference plots under development linked here
  5. Timeline for the next 10 days
    1. We have 9 days before we change targets, so we want to make sure we have enough statistics before we switch.
  6. Update on plugins

Wednesday, November 10

  1. Beam Update
  2. Production Run Update
    1. List of good runs is on the analysis shift page. Calibration can start for some subdetectors.
    2. Discussed the current progress regarding the run spreadsheet and the data quality spreadsheet
    3. Need to have a summary slides of the important plots to monitor for the data quality spreadsheet
  3. Physics on ρ0 channel of Run 90059
  4. Trigger Discussion
    1. 2 triggers: Main production:
      1. Bit 1 BCAL + FCAL + SC
      2. bit2= BCAL + FCAL with larger energy threshold
      3. bit 3 =PS
    2. FCAl and BCAL minimum ionizing (1)
    3. Start counter
    4. Will kill some physics, primakov production, excited nucleon or recoil on coherent only neutrals (2) in final state
    5. To account for this, add trigger with larger energy threshold to accept primakov style events
    6. Completelmentary, has larger energy deposition
    7. Primakov is z^2 hunting for small enhancement bumble in angle dis. Or eta, would be big peak with carbons.
      1. larger eta mass than pion
      2. larger the target the better, enhancement for primakov
      3. pi0 in hall b (primakov in carbon)
      4. more physics, sensitive to this
      5. 50 kHz with 90% live time
    8. Adding TPOL yesterday
      1. will use it somewhere eventually
    9. Ahead of schedule
      1. already have production data on he 4
      2. Took AMO that can used in
    10. For Flux,
      1. tagger and PS needs to be calibrated
      2. estimate is total number of PS triggers
      3. integral of PS spectrum
    11. Insert calorimeter into beam, count photons. In parallel, run PS, so you know acceptance, convulsion of PS converter (thickness and sigma) acceptance

Tuesday, November 09

  1. Radiator
  2. Timing Cuts
  3. Bo and Phoebe's input from the spreadsheets
    1. Discussed updates to Analysis shift and analysis shift webpage
  4. Possible take a look over the run files that we have, if they get ready

Monday, November 08